Write the equation of the line passing through (-1,7) and (-1,1)
Answer by checkley71(8403)

FIRST WE FIND THE SLOPE(Y2-Y1)/(X2-X1)
SLOPE=(1-7)/(-1+1)=-6/0 OR A VERTICLE LINE THROUGH THE POINT X=-1
THUS THE LINE EQUATION IS X=-1 ANSWER.